# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v1

@@ 3 Algorithm design
7:     Stopping test. Compute $t^{k+1}$ according to (22). Stop when $k\geq 1$ and $\max\{s_{k},\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\|,t^{k+1}\}\leq tol$.
8:     Penalty parameter update. Set $p_{k+1}$ according to (23).
9:  end for
10:  return  $(x^{k+1},y^{k+1})$
@@ 4.1 Preliminary results for the analysis
We begin by considering the following lemma, which provides an estimate for the distance between the approximation $\theta$ and the exact solution $\theta^{*}(x,y)$ of the proximal lower-level problem, based on the prox-gradient residual $\mathcal{G}(\theta,x,y)$ in the inexact criteria (18) and (21). This estimate is derived from the $1/\gamma$-strongly convexity and $1/\gamma+L_{f}$-smoothness of the component $f(x,\theta)+\|\theta-y\|^{2}/(2\gamma)$ with respect to $\theta$ in the objective of the proximal lower level problem, see, e.g., [53, Theorem 3.4 and 3.5] or [54, Proposition 2.2].
@@
$\|\theta-\theta^{*}(x,y)\|\leq C_{\eta}\left\|\theta-\operatorname{Prox}_{\eta\tilde{g}\left(x,\cdot\right)}\left(\theta-\eta\left(\nabla_{y}f\left(x,\theta\right)+(\theta-y)/{\gamma}\right)\right)\right\|,$
where $\theta^{*}(x,y):=\arg\min_{\theta\in Y}\{\varphi(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}\}$.
@@
Suppose Assumptions 2 and 3 hold. Let $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$ and $(\bar{x},\bar{y})\in X\times{\mathbb{R}}^{m}$. Then, for any $(x,y)\in X\times{\mathbb{R}}^{m}$, the following inequalities holds,
@@
According to Proposition 2, for any $\bar{y}\in{\mathbb{R}}^{m}$, $v_{\gamma}(x,\bar{y})$ is a $\rho_{f_{1}}+\rho_{g_{1}}$-weakly convex function with respect to $x$ on $X$, that is, $v_{\gamma}(x,\bar{y})+\frac{\rho_{f_{1}}+\rho_{g_{1}}}{2}\|x\|^{2}$ is convex on $X$. Then we can obtain (24) immediately. The second inequality (25) follows from Proposition 2 that for any $\bar{x}\in X$, $v_{\gamma}(\bar{x},y)$ is convex with respect to $y$. ∎
@@
Suppose Assumptions 1, 2 and 3 hold. Let $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, $\alpha_{k}\in[0,2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[0,2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $c_{\alpha},c_{\beta}$ are positive constants. Then, there exists $C_{\eta}>0$ such that the sequence $(x^{k},y^{k})$ generated by AGILS (1) satisfies the following inequality,
$\tilde{\psi}_{p_{k+1}}\left(x^{k+1},y^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)+\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}$ (26)
$-\frac{c_{\alpha}}{4}\left\|x^{k+1}-x^{k}\right\|^{2}-\frac{c_{\beta}}{4}\left\|y^{k+1}-y^{k}\right\|^{2}.$
@@
First, by the update rule for $y^{k+1}$ given in (15), we have
$g(x^{k},y^{k+1})+\langle d_{y}^{k},y^{k+1}-y^{k}\rangle+\frac{1}{\beta_{k}}\|y^{k+1}-y^{k}\|^{2}\leq g(x^{k},y^{k}).$ (27)
@@
$\leq$ $(\tilde{\psi}_{p_{k}}-g)(x^{k},y^{k})+\langle\nabla_{y}(\tilde{\psi}_{p_{k}}-g)(x^{k},y^{k}),y^{k+1}-y^{k}\rangle+\frac{L_{\psi_{y,k}}}{2}\|y^{k+1}-y^{k}\|^{2},$
@@
$\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)+\langle\nabla_{y}(\tilde{\psi}_{p_{k}}-g)(x^{k},y^{k})-d_{y}^{k},y^{k+1}-y^{k}\rangle$ (29)
$-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}\right)\left\|y^{k+1}-y^{k}\right\|^{2}.$
Using the formula of the gradient of $v_{\gamma}$ given in (9) and the construction of $d_{y}^{k}$ in (14), and 3 and the inexact condition (18), we have
$\left\|\nabla_{y}(\tilde{\psi}_{p_{k}}-g)(x^{k},y^{k})-d_{y}^{k}\right\|$ $=\|\nabla_{y}v_{\gamma}(x^{k},y^{k})+(y^{k}-\theta^{k})/\gamma\|$ (30)
$=\frac{1}{\gamma}\left\|\theta^{*}(x^{k},y^{k})-\theta^{k}\right\|\leq\frac{C_{\eta}}{\gamma}s_{k}.$
@@
$\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\leq\,\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-y^{k}\right\|^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}s_{k}^{2}.$ (31)
@@
$\left\|\nabla_{x}\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})-d_{x}^{k}\right\|$ $=\left\|\nabla_{x}\varphi(x^{k},\theta^{*}(x^{k},y^{k+1}))-\nabla_{x}\varphi(x^{k},\theta^{k+1/2})\right\|$ (35)
$\leq(L_{f_{x}}+L_{g_{2}})\left\|\theta^{*}(x^{k},y^{k+1})-\theta^{k+1/2}\right\|\leq(L_{f_{x}}+L_{g_{2}})C_{\eta}s_{k},$
where the last inequality follows from 3 and the inexact criterion (21). Combining the above inequality with (34) gives us
@@
$+\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}C_{\eta}^{2}s_{k}^{2}.$
Combining (31) and (36) yields the following decreasing property,
@@
$\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)+\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}C_{\eta}^{2}s_{k}^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}s_{k}^{2}$ (37)
$-\left(\frac{1}{\alpha_{k}}-\frac{L_{\psi_{x,k}}}{2}-\frac{c_{\alpha}}{4}\right)\left\|x^{k+1}-x^{k}\right\|^{2}-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-y^{k}\right\|^{2}.$

### v2

@@ 3 Algorithm design
7:     Calculate $\Delta_{k+1}:=\left\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\right\|$ and $t^{k+1}$ according to (21). Stop when $k\geq 1$ and $\max\{\Delta_{k+1},s_{k},t^{k+1}\}\leq tol.$
8:     Update penalty parameter and apply feasibility correction:
9:        Case 1: $\Delta_{k+1}\geq c_{p}\min\big{\{}1/p_{k},\,t^{k+1}\big{\}}$;      Set $p_{k+1}:=p_{k}$, $\tilde{y}^{k+1}:=y^{k+1}$, $\tilde{\theta}^{k+1}:=\theta^{k+1}$.
10:        Case 2: $\Delta_{k+1}<c_{p}\min\big{\{}1/p_{k},\,t^{k+1}\big{\}},\,\|y^{k+1}-\theta^{k+1}\|\leq c_{y}\gamma/p_{k};$      Set $p_{k+1}:=p_{k}+\varrho_{p}$, $\tilde{y}^{k+1}:=y^{k+1}$, $\tilde{\theta}^{k+1}:=\theta^{k+1}$.
11:        Case 3: $\Delta_{k+1}<c_{p}\min\big{\{}1/p_{k},\,t^{k+1}\big{\}},\,\|y^{k+1}-\theta^{k+1}\|>c_{y}\gamma/p_{k};$      Apply the feasibility correction procedure (Algorithm 2).
12:  end for
13:  return  $(x^{k+1},\tilde{y}^{k+1})$
Algorithm 2 Feasibility Correction Procedure
1:  Input: Current iterates $x^{k+1},y^{k+1},\theta^{k+1}$, $\Delta_{k+1}$, proximal parameter $\gamma$, penalty parameters $p_{k}$, $\varrho_{p}$, inexact parameters $\eta$, $s_{k}$, $\tau_{k}$, constant $c_{\tilde{y}}$.
2:  Compute a candidate $\tilde{y}^{k+1}$ by inexactly solving the lower-level problem (2) with $x=x^{k+1}$ to satisfy (24).
3:  Find an inexact solution $\tilde{\theta}^{k+1}$ that satisfies the inexactness criterion (26).
4:  if the descent condition (25) holds then
5:     Accept the candidate: Set $p_{k+1}:=p_{k}$.
6:  else
7:     Reject the candidate: Set $\tilde{y}^{k+1}:=y^{k+1}$, $\tilde{\theta}^{k+1}:=\theta^{k+1}$ and $p_{k+1}:=p_{k}+\varrho_{p}$.
8:  end if
9:  return  $(\tilde{y}^{k+1},\tilde{\theta}^{k+1},p_{k+1})$
@@ 4.1 Preliminary results for the analysis
We begin by considering the following lemma, which provides an estimate for the distance between the approximation $\theta$ and the exact solution $\theta^{*}_{\gamma}(x,y)$ of the proximal lower-level problem, based on the prox-gradient residual $\mathcal{G}(\theta,x,y)$ in the inexact criteria (17) and (20). This estimate is derived from the $1/\gamma$-strong convexity and $(1/\gamma+L_{f})$-smoothness of the component $f(x,\theta)+\|\theta-y\|^{2}/(2\gamma)$ with respect to $\theta$ in the objective of the proximal lower level problem, see, e.g., [23, Theorem 3.4 and 3.5] or [48, Proposition 2.2].
@@
$\|\theta-\theta^{*}_{\gamma}(x,y)\|\leq C_{\eta}\left\|\theta-\operatorname{Prox}_{\eta\tilde{g}\left(x,\cdot\right)}\left(\theta-\eta\left(\nabla_{y}f\left(x,\theta\right)+(\theta-y)/{\gamma}\right)\right)\right\|,$
where $\theta^{*}_{\gamma}(x,y):=\arg\min_{\theta\in Y}\{\varphi(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}\}$.
@@
Suppose Assumptions 2 and 3 hold. Let $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$ and $(\bar{x},\bar{y})\in X\times{\mathbb{R}}^{m}$. Then, for any $(x,y)\in X\times{\mathbb{R}}^{m}$, the following inequalities hold,
@@
According to Proposition 2, for any $\bar{y}\in{\mathbb{R}}^{m}$, $v_{\gamma}(x,\bar{y})$ is a $(\rho_{f_{1}}+\rho_{g_{1}})$-weakly convex function with respect to $x$ on $X$, that is, $v_{\gamma}(x,\bar{y})+\frac{\rho_{f_{1}}+\rho_{g_{1}}}{2}\|x\|^{2}$ is convex on $X$. Then we can obtain (27) immediately. The second inequality (28) follows from Proposition 2 that for any $\bar{x}\in X$, $v_{\gamma}(\bar{x},y)$ is convex with respect to $y$. ∎
@@
Before establishing the main decreasing property, we first demonstrate that the descent condition (25), imposed by the feasibility correction procedure (Algorithm 2), ensures a decrease in $\tilde{\psi}_{p_{k}}$ when moving from $y^{k+1}$ to the corrected iterate $\tilde{y}^{k+1}$.
@@
Suppose Assumptions 2 and 3 hold. Then, there exists $C_{\gamma}>0$, such that the sequence $\{(x^{k},y^{k},\tilde{y}^{k})\}$ generated by AGILS (Algorithm 1) satisfies
$\tilde{\psi}_{p_{k}}(x^{k+1},\tilde{y}^{k+1})\leq\tilde{\psi}_{p_{k}}(x^{k+1},y^{k+1})+C_{\gamma}\zeta_{k+1}^{2},$ (29)
where $\zeta_{k}:=\max\{s_{k},(\prod_{i=0}^{k}\tau_{i})\mathcal{G}(\theta^{0},x^{0},y^{0})\}$.
@@
At a given iteration $k$, if $\tilde{y}^{k+1}=y^{k+1}$, the inequality holds trivially.
Now, consider the case where $\tilde{y}^{k+1}\neq y^{k+1}$. This implies that the feasibility correction procedure (Algorithm 2) is applied at iteration $k$. Accordingly, the corrected iterate $\tilde{y}^{k+1}$ satisfies condition (25), and its corresponding $\tilde{\theta}^{k+1}$ satisfies condition (26). Condition (26) implies that $\mathcal{G}(\tilde{\theta}^{k+1},x^{k+1},\tilde{y}^{k+1})\leq\zeta_{k+1}$.
Assumptions imply the $1/\gamma$-strong convexity and $(1/\gamma+L_{f})$-smoothness of the component $f(x,\theta)+\|\theta-y\|^{2}/(2\gamma)$ with respect to $\theta$. Recalling results from [37, Appendix F] and [7, Theorem 10.9], there exists $C_{\gamma}>0$ such that for any $\theta$,
$v_{\gamma}(x,y)=\inf_{\theta\in Y}\{\varphi(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}\}\geq\varphi(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}-C_{\gamma}\mathcal{G}(\theta,x,y)^{2}.$
Apply this to $x^{k+1}$, $\tilde{y}^{k+1}$ and $\tilde{\theta}^{k+1}$,
$-v_{\gamma}(x^{k+1},\tilde{y}^{k+1})$ $\leq-\varphi(x^{k+1},\tilde{\theta}^{k+1})-\frac{1}{2\gamma}\|\tilde{\theta}^{k+1}-\tilde{y}^{k+1}\|^{2}+C_{\gamma}\mathcal{G}(\tilde{\theta}^{k+1},x^{k+1},\tilde{y}^{k+1})^{2}$
$\leq-\varphi(x^{k+1},\tilde{\theta}^{k+1})-\frac{1}{2\gamma}\|\tilde{\theta}^{k+1}-\tilde{y}^{k+1}\|^{2}+C_{\gamma}\zeta_{k+1}^{2}$
By condition (25), which $\tilde{y}^{k+1}$ satisfies, we have
$\quad\tilde{\psi}_{p_{k}}(x^{k+1},\tilde{y}^{k+1})-\tilde{\psi}_{p_{k}}(x^{k+1},y^{k+1})$
$\leq-v_{\gamma}(x^{k+1},\tilde{y}^{k+1})+\left(\varphi(x^{k+1},\tilde{\theta}^{k+1})+\frac{1}{2\gamma}\|\tilde{\theta}^{k+1}-\tilde{y}^{k+1}\|^{2}\right)$
$\quad+v_{\gamma}(x^{k+1},{y}^{k+1})-\left(\varphi(x^{k+1},\theta^{k+1})+\frac{1}{2\gamma}\|\theta^{k+1}-y^{k+1}\|^{2}\right)$
$\leq C_{\gamma}\zeta_{k+1}^{2},$
where the last inequality follows from the definition of $v_{\gamma}(x^{k+1},{y}^{k+1})$ as an infimum $\inf_{\theta\in Y}\{\varphi(x^{k+1},{\theta})+\frac{1}{2\gamma}\|{\theta}-y^{k+1}\|^{2}\}$. Thus, the conclusion follows. ∎
Next, we establish a key recursive inequality for $\tilde{\psi}_{p_{k}}$, demonstrating a sufficient decrease property for the iterates generated by Algorithm 1.
Lemma 5.
Suppose Assumptions 1, 2 and 3 hold. Let $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, $\alpha_{k}\in[0,2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[0,2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $c_{\alpha},c_{\beta}$ are positive constants. Then, there exists $C_{\eta},C_{\gamma}>0$ such that the sequence $\{(x^{k},y^{k},\tilde{y}^{k})\}$ generated by AGILS (Algorithm 1) satisfies the following inequality,
$\tilde{\psi}_{p_{k+1}}\left(x^{k+1},\tilde{y}^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},\tilde{y}^{k}\right)+\left(\frac{C_{\eta}^{2}}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}\right)\zeta_{k}^{2}+C_{\gamma}\zeta_{k+1}^{2}$ (30)
$-\frac{c_{\alpha}}{4}\left\|x^{k+1}-x^{k}\right\|^{2}-\frac{c_{\beta}}{4}\left\|y^{k+1}-\tilde{y}^{k}\right\|^{2},$
where $\zeta_{k}:=\max\{s_{k},(\prod_{i=0}^{k}\tau_{i})\mathcal{G}(\theta^{0},x^{0},y^{0})\}$ and $\{\zeta_{k}\}$ is square summable.
Proof.
First, the update rule for $y^{k+1}$ given in (14) defines it as the minimizer of the function $g(x^{k},y)+\langle d_{y}^{k},y-\tilde{y}^{k}\rangle+\frac{1}{2\beta_{k}}\left\|y-\tilde{y}^{k}\right\|^{2}$. This function is $1/\beta_{k}$-strongly convex in $y$. By applying [7, Theorem 6.39], we obtain
$g(x^{k},y^{k+1})+\langle d_{y}^{k},y^{k+1}-\tilde{y}^{k}\rangle+\frac{1}{\beta_{k}}\|y^{k+1}-\tilde{y}^{k}\|^{2}\leq g(x^{k},\tilde{y}^{k}).$ (31)
@@
$\leq$ $(\tilde{\psi}_{p_{k}}-g)(x^{k},\tilde{y}^{k})+\langle\nabla_{y}(\tilde{\psi}_{p_{k}}-g)(x^{k},\tilde{y}^{k}),y^{k+1}-\tilde{y}^{k}\rangle+\frac{L_{\psi_{y,k}}}{2}\|y^{k+1}-\tilde{y}^{k}\|^{2},$
@@
$\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},\tilde{y}^{k}\right)+\langle\nabla_{y}(\tilde{\psi}_{p_{k}}-g)(x^{k},\tilde{y}^{k})-d_{y}^{k},y^{k+1}-\tilde{y}^{k}\rangle$ (33)
$-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}\right)\left\|y^{k+1}-\tilde{y}^{k}\right\|^{2}.$
Using the formula of $\nabla_{y}v_{\gamma}$ in (9), and the construction of $d_{y}^{k}$ in (13), we have
$\left\|\nabla_{y}(\tilde{\psi}_{p_{k}}-g)(x^{k},\tilde{y}^{k})-d_{y}^{k}\right\|$ $=\|\nabla_{y}v_{\gamma}(x^{k},\tilde{y}^{k})+(\tilde{y}^{k}-\theta^{k})/\gamma\|$ (34)
$=\frac{1}{\gamma}\left\|\theta^{*}_{\gamma}(x^{k},\tilde{y}^{k})-\theta^{k}\right\|$
$\leq\frac{C_{\eta}}{\gamma}\mathcal{G}({\theta}^{k},x^{k},\tilde{y}^{k})\leq\frac{C_{\eta}}{\gamma}\zeta_{k},$
where the first inequality follows from Lemma 2 and the second inequality is a consequence of the inexactness condition (20).
@@
$\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\leq\,\tilde{\psi}_{p_{k}}\left(x^{k},\tilde{y}^{k}\right)-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-\tilde{y}^{k}\right\|^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}\zeta_{k}^{2}.$ (35)
@@
$\left\|\nabla_{x}\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})-d_{x}^{k}\right\|$ $=\left\|\nabla_{x}\varphi(x^{k},\theta^{*}_{\gamma}(x^{k},y^{k+1}))-\nabla_{x}\varphi(x^{k},\theta^{k+1/2})\right\|$ (39)
$\leq(L_{f_{x}}+L_{g_{2}})\left\|\theta^{*}_{\gamma}(x^{k},y^{k+1})-\theta^{k+1/2}\right\|\leq(L_{f_{x}}+L_{g_{2}})C_{\eta}\zeta_{k},$
where the last inequality follows from Lemma 2 and the inexact criterion (17). Combining the above inequality with (38) gives us
@@
$+\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}C_{\eta}^{2}\zeta_{k}^{2}.$
Combining (35), (40), and then (29) established in Lemma 4, we obtain
@@
$\tilde{\psi}_{p_{k}}\left(x^{k+1},\tilde{y}^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},\tilde{y}^{k}\right)+\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}C_{\eta}^{2}\zeta_{k}^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}\zeta_{k}^{2}+C_{\gamma}\zeta_{k+1}^{2}$ (41)
$-\left(\frac{1}{\alpha_{k}}-\frac{L_{\psi_{x,k}}}{2}-\frac{c_{\alpha}}{4}\right)\left\|x^{k+1}-x^{k}\right\|^{2}-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-\tilde{y}^{k}\right\|^{2}.$